what is the answer to 1/10+1/6+3/5
Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

what is the answer to 1/10+1/6+3/5
---
least common denominator: 30
---
Convert each denominator to 30:
3/30 + 5/30 + 18/30
----
Combine the numerators over the lcd:
(3+5 + 18)/30
---
= (26/30)
---
= 13/15
================
Cheers,
Stan H.
================

Answer by rfer(16322)

6/60+10/60+36/60=52/60=13/15